OpenStudy (anonymous):

fSolve each equation. Leave the answer as an improper fraction. 3x2 + 2 = 11 - x2

OpenStudy (uri):

3x^2+x^2=11-2 Solve further? @dfrazier2

OpenStudy (texaschic101):

3x^2 + 2 = 11 - x^2 --- add x^2 to both sides 3x^2 + x^2 + 2 = 11 --- subtract 2 from each side 3x^2 + x^2 = 11 - 2 --- combine like terms 4x^2 = 9 --- divide both sides by 4 x^2 = 9/4 -- take sq rt of each side x = sq rt 9/4 x = sq rt 2 1/4 or 2.25 x = 1.5

OpenStudy (texaschic101):

oops...leave as improper fraction.... x = sq rt 9/4 ??

OpenStudy (anonymous):

Thankz

OpenStudy (texaschic101):

no problem :)